Suppose a firm produces according to the production function Q = 2L0.6K0.2, and faces wage rate

₵10, a rental cost of capital ₵5, and sells output at a price of ₵20.

a. Obtain and expression for the factor demand functions.

b. Compute the profit-maximizing factor demands for capital and labour.


1
Expert's answer
2021-10-15T14:20:19-0400

Solution:

a.). Input demand for labor: L = F(w,r,p)

Input demand for Capital: K = F(w,r,p)

TC = wL + rK = 10L + 5K

TR = P "\times" Q = 20(2L0.6K0.2)

Profit = PQ – wL – rK = 20(2L0.6K0.2) - 10L - 5K

 

b.). Profit = PQ – wL – rK = 20(2L0.6K0.2) - 10L - 5K

"\frac{\partial \pi } {\partial L} =" 24L-0.4K0.2 – 10 = 0        (a)

 

"\frac{\partial \pi } {\partial K} =" = 8L0.6K-0.8 – 5 = 0           (b)


Solve for equation (a) for L:

24L-0.4K0.2 = 10

L = 8.9K0.5

Plug into equation (b):

8(8.9K0.5)0.6K-0.8 = 5

29.7K0.3K-0.8 = 5

29.7K-0.5 = 5

K = 35.28


Substitute to derive L:

L = 8.9K0.5

L = 8.9(35.280.5) = 52.86

L = 52.86


Q = 2L0.6K0.2

Q = 2(52.860.6) (35.280.2) = 21.62 "\times" 2.04 = 44.10

Q = 44


Derive Profit:

Profit = 20(44) – 10(52.86) – 5(35.28)

Profit = 880 – 528.6 – 176.4

Profit = 175